A sample was done, collecting the data below. Calculate the standard deviation, to one decimal place.

x
22
1
2
27
16
Answer #1

# standard devinson - Zcxi-7,2 n-l here ti Exi = 22+1+2+27+ 16 = 13.6 s - 22-13.6) 3+ (1-13.6)2+(2-13.602 + (27-13.632 + ( 16